OpenStudy (anonymous):

derivative of sqrt(x^2 + (sinX)^2)

OpenStudy (fibonaccichick666):

\[\sqrt{x^2 + (sinX)^2}\] this?

OpenStudy (anonymous):

I looked the answer up, and it's wrong, just not sure why...

OpenStudy (fibonaccichick666):

looks like you forgot to cancel the 2

OpenStudy (fibonaccichick666):

other than that at first glance I don't see anything else wrong. I do recommend using substitutions for the chain rule though

OpenStudy (anonymous):

what are substitutions, do you ever start from the innermost and work outward?

OpenStudy (anonymous):

wolfram says I'm wrong http://www.wolframalpha.com/input/?i=derivative+of+sqrt%28x%5E2+%2B+%28sinX%29%5E2%29

OpenStudy (fibonaccichick666):

ignore my last, ok let's do it then

OpenStudy (anonymous):

have to run, can you work through the steps for me?

OpenStudy (fibonaccichick666):

\[f(x)=\sqrt{x^2 + (sinx)^2}\]

OpenStudy (fibonaccichick666):

this takes 2 minutes come on

OpenStudy (anonymous):

literally have three minutes

OpenStudy (fibonaccichick666):

let u=x^2+sin^2x

OpenStudy (fibonaccichick666):

du=?

OpenStudy (fibonaccichick666):

then you have \[f(x)=u^\frac{1}{2}\]

OpenStudy (fibonaccichick666):

f'(u)=?

OpenStudy (fibonaccichick666):

then backsub

OpenStudy (fibonaccichick666):

don't forget make \(r=sinx\) then dr=cosx

OpenStudy (fibonaccichick666):

so you get du=?...dr

OpenStudy (fibonaccichick666):

then that's it

OpenStudy (fibonaccichick666):

when you put in two different sized x's it did the partial with respect to x

OpenStudy (fibonaccichick666):

You learn about those in calc III